05 September 2011 UT

Clouds, high humidity and rain.

Check out basic functionality with dome closed.

  • Some of the test OBs used tonight are in /home/okuhn/Calib_OBs/Test20110905/.

  • LBC Red exposure times. Previously, to certain commanded exposure times, a 1.3 sec offset was added to the reported time in the headers (reported as EXPTIME and TEXPTIME). Fernando reported that this was due to an electronics problem and was fixed. Took a series of exposures at some of the problematic exposure times and verified that the EXPTIME, TEXPTIME were what were expected, and no 1.3 offset was added.

  • Ran an OB which slewed, moved filters and dithered. Started with a non-zero dither.
    • non-zero first dither worked as expected (had been fixed earlier).
    • when slewing and moving rotator (from PA=+100 to PA=-100), noted that on the first pair of images, the exposure time reported in the User Interface was PA=-173 deg, not PA=-100. Before the second pair of images was started, the rotator moved so that the PA for these would be -100 deg. Header reads PA=-100, but the header is written as the image was being saved. Was the PA really not correct for the first pair of images (I suspect it was not correct, but this must be checked on-sky). IT 3595.
    • Noticed lbcbtec images look noisy. Lights were off in dome and all were close to bias levels. With lights on, these look OK. =_Not really --- during observing, the light levels would be rather low, and for ~500 ADU, the stddev in a 5x5 pixel box was ~25-30 ADU. For comparison, stddev ~ 5 ADU for Red. This could affect finding and measuring pupils on the tech chip. IT #3596 submitted.

  • Ran the Test OB of Matthias Mueller to measure overheads. This OB is set to read only 2016 rows, so window is [2304:2016]. It takes 3 exposures at z-SLOAN, 3 at V-BESSEL, 3 at z-SLOAN... The time between exposures in the same filter is 24-26 seconds and the time between these when there is a filter change is 28 seconds (z-SLOAN -> V-BESSEL) and 27 seconds (V-BESSEL -> z-SLOAN). I ran the OB twice and get the same differences for each readout and move.

  • Slewed again to test dithering, rotator, lbcbtec images with lights on. Filter Wheel #1 stop return error (out:) Reported filter error in IT 3594.

  • Ran OB (/home/okuhn/Calib_OBs/DitherTest.ob).

  • Woke Fernando to look at the high noise on lbcbtec images. Turned off telescope tracking and HBS but still see high noise. But in the meantime, we see also a strange pattern on a science image, with vertical lines through chips 1 & 3 and a series of bars along the left edges of chips 1,3 and 4. See lbcb.20110905.103238.fits for an example. This diminishes over the next lbcb images.
    • Fernando noted that the Blue Camera & Trackers electronics temperatures were high, both ~70 C (for comparison, for the red side these are ~55 C). We turned off/on the camera & trackers. But even when Blue Trackers electronics temperature was 55 C, there was noise pattern, see lbcbtec.20110905.11*_1/2.fits
    • Fernando in dome checking cabling. He disconnected ground, but we still see noise - no change.

  • In Blue Science images, we see strange patterns. Two types: 1/ "wedge/triangle" shapes where counts are slightly higher. 2/ 512 row segments on left sides of chips, and vertical bar, creating a sort of step function along X direction. First time I noticed these was when taking 45 sec images (so tech chip was triggered). Examples are (of 1) lbcb.20110905.064807.fits, and (of 2) lbcb.20110905.065456.fits.

  • One image did not transfer to the archive - only the header was written to /newdata so far. This image is lbcb.20110905.082807.fits

  • Looking at last night's data, I see this pattern appearing in a 1 sec image lbcb.20110904.071558.fits. This pattern appears in the biases also. It may arise because the hard flushing had been removed from the sequence in some instances. Andrea is editing the LBC control code.

-- OlgaKuhn - 05 Sep 2011
Topic revision: r2 - 05 Sep 2011, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback